OpenStudy (anonymous):

The driver of a car slams on the brakes when he sees a tree blocking the road. The car slows uniformly with acceleration of −4.10 m/s2 for 3.30 s, making straight skid marks 60.0 m long, all the way to the tree. With what speed does the car then strike the tree?

OpenStudy (anonymous):

Applying \[v ^{2}=u^{2}+2as\] we have v^2 = u^2 - 8.2 x 60. Then applying \[s = ut + \frac{ 1 }{ 2}at ^{2}\] we see that 60 = 3.30u-2.05x3.3^2 and so the initial speed was 24.94 m/s. Putting this into the first equation, we find that v = 11.42 m/s.

OpenStudy (anonymous):

In conclusion, the driver is in hospital with major injuries.